Wayne charges the following for repairing washing machines:

£28 call-out charge + £16 for each half-hour he spends on the repair

What will the total cost be for a repair which takes one and a half hours?


1
Expert's answer
2022-01-17T17:53:32-0500

Out-charge= C

charges per hour= x

Total cost= y

In one and a half hours= 3half hours

"y=mx+c"

"=16x+28"

"=(16*3)+28"

"=48+28"

"=76"

The total repair cost is "\epsilon76"